# Fast Transaction Scheduling in Blockchain Sharding

### v1

@@
Abstract.
Sharding is a promising technique for addressing the scalability issues of blockchain. It divides the $n$ participating nodes into $s$ disjoint groups called shards, where each shard processes transactions in parallel. The accounts (shared objects) are also distributed among the $s$ shards. We investigate scheduling algorithms for the blockchain sharding systems, where each transaction resides in a shard of the communication graph and attempts to access accounts at possibly remote shards. We examine batch scheduling problems on the shard graph $G_{s}$, where given a set of transactions, we aim to find efficient schedules to execute them as fast as possible. First, we present a centralized scheduler where one of the shards has global knowledge of transactions to be processed. For general graphs, where the transaction and its accessing objects are arbitrarily far from each other with a maximum distance $d$, the centralized scheduler provides $O(kd)$ approximation to the optimal schedule, where $k$ is the maximum number of shards each transaction accesses. Consequently, for a Clique graph where shards are at a unit distance from each other, we obtain $O(k)$ approximation to the optimal schedule. We also get $O(k\log s)$ approximation for Hypercube, Butterfly, and $g$-dimensional Grid, where $g=O(\log s)$. Next, we provide a centralized scheduler with a bucketing approach that offers improved bounds for special cases. For the General graph where each transaction chooses $k$ objects randomly we give $O(k\log D\cdot(k+\log s))$ approximation, and for the Line graph we provide $O(k\sqrt{d}\log D)$ approximation to the optimal schedule, where $D$ is the diameter of shard graph $G_{s}$. Finally, we provide a distributed scheduler where shards do not require global transaction information. We achieve this by using a hierarchical clustering of the shards and using the centralized scheduler in each cluster. We show that the distributed scheduler has a competitive ratio of $O(\mathcal{A_{\mathcal{CS}}}\log^{2}s)$, where $\mathcal{A_{\mathcal{CS}}}$ is the approximation ratio of the centralized scheduler. To our knowledge, we are the first to give provably fast transaction scheduling algorithms for blockchain sharding systems.
@@ Abstract.
ccs: Computing methodologies Distributed algorithmsccs: Theory of computation Scheduling algorithms.
1. Introduction
@@
Blockchains are known for their special features such as fault tolerance, transparency, non-repudiation, immutability, and security (Sankar et al., 2017). Thus, blockchains have been used in various domains, such as cryptocurrency (Nakamoto, 2009; Buterin et al., 2014), healthcare (McGhin et al., 2019), digital forensics (Troyer et al., 2021; Akbarfam et al., 2023b, a), supply chain management (Azzi et al., 2019), edge and scientific computing (Al-Mamun et al., 2022, 2021).
Blockchain technology is also gaining interest from banks and financial institutions because of its transformational power in the financial industry (Androulaki et al., 2020; Javaid et al., 2022; Alamsyah et al., 2024; Weerawarna et al., 2023). This innovation is expected to reshape financial services, revolutionize banking operations, and address dynamic industry demands without sacrificing privacy (Baum et al., 2023; Weerawarna et al., 2023). Implementing blockchain can offer competitive benefits to traditional banking systems by enhancing transaction privacy and security at a minimum cost (Cucari et al., 2022). Traditional financial institutions, tech firms, and startups are also coordinating with financial technology to offer affordable and easy to use financial services (Canaday, 2017). However, the primary drawback of blockchain is its scalability and its performance, which is dependent on the throughput of the transactions (Cocco et al., 2017; Chen and Bellavitis, 2020). Thus, there is a research gap to address the scalability issues of blockchain.
@@
To append a new block in the blockchain network, each participating node needs to reach a consensus, which is a time and energy-consuming process (Adhikari and Busch, 2023; Adhikari et al., 2024a, b). Moreover, each node is required to process and store all transactions, which leads to scalability issues in the blockchain system. To improve scalability and performance, sharding protocols have been proposed, such as Elastico (Luu et al., 2016), OmniLedger (Kokoris-Kogias et al., 2018), RapidChain (Zamani et al., 2018), Sharper (Amiri et al., 2021), and ByShard (Hellings and Sadoghi, 2021). However, these sharding protocols do not provide a formal analysis of how we can schedule the transaction as fast as possible. Sharding divides participating nodes into smaller groups called shards that allow the parallel processing of transactions. In the sharded blockchain, independent transactions are processed and committed in multiple shards concurrently, which improves the blockchain system’s throughput.
@@
We consider a blockchain system consisting of $n$ nodes, which are further divided into $s$ shards, and each shard consists of a set of nodes. Shards are connected in a graph network $G_{s}$ with a diameter $D$, and each shard holds a subset of the objects (accounts). We assume that transactions are distributed across the shards, and each transaction accesses at most $k$ accounts. A transaction $T$ initially is in one of the shards; we called it as home shard for $T$. Similar to other sharding systems (Hellings and Sadoghi, 2021; Adhikari and Busch, 2023; Adhikari et al., 2024a, b), each transaction $T$ is split into subtransactions, where each subtransaction accesses an account. A subtransaction of $T$ is sent to the destination shard that holds the respective account. The maximum distance between the home shard and the destination shards in the shard graph is at most $d\leq D$. In other words, the transaction and its accessing shards (objects) are arbitrarily far from each other with a maximum distance of $d$. All home shards process transactions concurrently. A problem occurs when conflicting transactions try to access the same account simultaneously. In such a case, the conflict prohibits the transactions to be committed concurrently and forces them to serialize (Adhikari et al., 2024a, b). Our proposed scheduling algorithms coordinate the home shards and destination shards to process the transactions (and respective subtransactions) in a conflict-free manner. Each destination shard maintains a local blockchain of the subtransactions that are sent to it. The global blockchain can be constructed by taking the union of the local blockchains at the shards (Adhikari and Busch, 2023).
@@
We consider batch problem instances where a (batch) set of transactions ${\mathcal{T}}$ initially reside at shards and need to be processed. We study the shard model in different graph topologies such as General graph, Clique (Complete) graph, Hypercube, Butterfly, Grid, and Line graphs and provide the analysis for execution time. The scheduler determines the time step for each transaction $T_{i}\in{\mathcal{T}}$ to process and commit. The execution of our scheduling algorithm is synchronous, where the timeline is divided into time steps referred to as rounds. Similar to the work in (Adhikari et al., 2024a, b), we consider that the duration of a round is sufficient to allow the execution of the PBFT consensus algorithm (Castro et al., 1999) in each shard. A round is also the time to send a message between shards in a unit distance. The main goal of a scheduling algorithm is to efficiently and fairly process all transactions while minimizing the total execution time (makespan) until all transactions are either committed or aborted. Unlike previous sharding approaches (Hellings and Sadoghi, 2021; Kokoris-Kogias et al., 2018; Zamani et al., 2018), our scheduling algorithms do not require locking mechanisms for concurrency control. Hence, we can also say our algorithms are lock-free.
Shards Connected as Approximation to Optimal Schedule of
@@
General Graph $O(kd)$ $O(kd\log^{2}s)$
Hypercube, Butterfly, and
$g$-dimensional Grid Graph $O(k\log s)$ $O(k\log^{3}s)$
General Graph where,
$k$ is chosen as random $O(k\log D\cdot(k+\log s))$ $O(k\log D\cdot(k+\log s)\log^{2}s)$
Line Graph $O(k\sqrt{d}\log D)$ $O(k\sqrt{d}\log D\cdot\log^{2}s)$
@@
Table 1. Summary of the proposed centralized and distributed scheduler approximation to the optimal schedule. The used notations are as follows: $s$ represents the total number of shards, $k$ denotes the maximum number of shards (objects) accessed by each transaction, $d$ denotes the worst distance between any transaction (home shard) and its accessing objects (destination shard), $D$ denotes the diameter of shard graph $G_{s}$.
1.1. Contributions and Techniques
@@
To our knowledge, we give the first provably efficient batch transaction scheduling algorithms for blockchain sharding systems. We provide a summary of contributions in Table 1, which are also described as follows:
@@
  • Centralized Scheduler: First, we provide a Centralized scheduler where one of the shards has global knowledge of the transactions. Consider a general graph $G_{s}$ with diameter $D$. Suppose that each transaction accesses at most $k$ objects, where the distance between a transaction and its accessing objects is at most $d\leq D$. The centralized scheduler provides an $O(kd)$ approximation to the optimal schedule. Hence, when $G_{s}$ is a Clique graph, where shards are at a unit distance from each other (thus, $d=1$), we obtain an $O(k)$ approximation to the optimal schedule. Moreover, we obtain an $O(k\log s)$ approximation for the Hypercube, Butterfly, and $g$-dimensional Grid where $g=O(\log s)$, since in all these cases $d=O(\log s)$.
    • Technique: Our scheduling algorithm consists of two main phases. In the first phase, the scheduler determines the order of execution of the transactions by coloring transactions that need to be processed using a conflict graph. In the second phase, our algorithm confirms and commits each transaction according to the schedule determined by the first phase. In the centralized scheduler, we consider a set of transactions ${\mathcal{T}}$ that are distributed in a shard graph $G_{s}$, and one of the shard as leader shard are given to the algorithm as input. Then, the scheduler provides the schedule and processes transactions ${\mathcal{T}}$. We assume that the leader shard has global knowledge of the transactions that need to proceed. Suppose $G_{\mathcal{T}}$ is the transactions conflict graph where each node represents a transaction, and an edge between two nodes corresponds to a conflict between them. In the first phase, the leader shard schedules the transactions using a greedy coloring of $G_{\mathcal{T}}$, which assigns execution times to each transaction based on the coloring of $G_{\mathcal{T}}$. A valid coloring of $G_{\mathcal{T}}$ assigns a unique positive integer to each transaction such that two adjacent transactions (conflicting transactions) receive different colors. The colors correspond to the distinct time steps (we also called round) where the respective transactions are processed for commit. After coloring, the leader shard shares the color information with each home shard. Then, in the second phase, each home shard processes its transactions to commit in the destination shards. We described the proposed centralized scheduling algorithm in Section 4.
@@
  • Centralized Scheduler with Buckets: Next, we give the centralized scheduler with a bucketing approach. Similar to a centralized scheduler, we consider a shard graph $G_{s}$ with diameter $D$, and the maximum distance between the home shard of the transaction and its accessing objects is at most $d\leq D$. We divide the transactions into buckets $B_{0},B_{1},\ldots$, according to the maximum distance between the home shard of the transaction and its accessing objects (i.e., destination shards). $B_{i}$ consists of the transactions whose maximum distance between the home shard and its destination shard is in range $[2^{i},2^{i+1})$. This approach offers improved bounds for special cases: for the General graph where the $k$ objects are chosen randomly per transaction we give $O(k\log D\cdot(k+\log s))$ approximation, and for the Line graph, we provide $O(k\sqrt{d}\log D)$ approximation to the optimal schedule.
    • Technique: Each home shard of transaction $T_{i}\in{\mathcal{T}}$ discovers the position of its accessing objects (destination shards) and sends that transaction information to the respective bucket $B_{i}$. Bucket $B_{i}$ denotes the set of transactions whose distance between the home shard and the accessing objects (destination shards) is in the range $[2^{i},2^{i+1})$; $i\in\{0,\dots,\lfloor\log D\rfloor\}$ where $D$ is the diameter of shards graph $G_{s}$. We assume that each bucket $B_{i}$ has its own leader shard $\ell_{i}$, which is known to all shards in $G_{s}$. First, each transaction determines its bucket, and then we run the centralized scheduler for each bucket. Details of the algorithm is provided in Section 5.
@@
  • Distributed Scheduler: Finally, we give a distributed scheduler where shards are not aware of global transaction information. We use a clustering approach for the shard network by hierarchical clustering (Gupta et al., 2006) so that independent transactions can be scheduled and committed concurrently. The basic idea is to simulate the centralized scheduler within each cluster of the hierarchy to make a distributed scheduler. We show that the distributed scheduler has a competitive ratio of $O(\mathcal{A_{\mathcal{CS}}}\log^{2}s)$, where $\mathcal{A_{\mathcal{CS}}}$ is the approximation ratio of the centralized scheduler.
    • Technique: We consider that our distributed scheduling algorithm uses a hierarchical decomposition of shard graph $G_{s}$ which is known to all the shards and calculated before the algorithm starts. This shard clustering (graph decomposition) is based on the clustering techniques in (Gupta et al., 2006) and which were later used in (Sharma and Busch, 2014; Busch et al., 2022; Adhikari et al., 2024a, b). We divide the shard graph $G_{s}$ into the hierarchy of clusters with layers and sublayers. The home cluster for each transaction $T_{i}$ is defined as follows: suppose $S_{i}$ is the home shard of $T_{i}$, and $z$ is the maximum distance from $S_{i}$ to the destination shards that will be accessed by $T_{i}$; the home cluster of $T_{i}$ is the lowest-layer (and lowest sub-layer) cluster in the hierarchy that contains the $z$-neighborhood of $S_{i}$. Each home cluster has one dedicated leader shard, which will handle all the transactions that have their home shard in that cluster (i.e., transaction information is sent from the home shard to the cluster leader shard for processing). After that, each cluster leader shard schedules the transaction of its cluster by assigning color using a greedy coloring algorithm (similar to the centralized scheduler). To achieve the benefit of locality and reduce the transaction latency, the lower cluster layer gets priority in processing its transactions. Within each cluster, we simulate and run a centralized scheduler algorithm to schedule transactions. A detailed description of our distributed scheduler is in Section 6.
@@
Paper Organization: The rest of this paper is structured as follows: Section 2 provides related works. Section 3 describes the preliminaries for this study and the sharding model. Section 4 presents a centralized scheduling algorithm. In Section 5, we provide a centralized scheduler with the bucketing approach. Section 6 generalizes the techniques to a fully distributed setting. Finally, we give our conclusions in Section 7.
2. Related Work
@@
Extensive research has been proposed to address the scalability issues of blockchain in the consensus layer (Jalalzai and Busch, 2018; Jalalzai et al., 2019, 2021). While these proposed algorithms and protocols have made some improvements in scalability, the system’s performance still degrades as the network size increases. To tackle the scalability issue of blockchain, various sharding protocols (Luu et al., 2016; Kokoris-Kogias et al., 2018; Zamani et al., 2018; Hellings and Sadoghi, 2021; Dang et al., 2019; Amiri et al., 2021; Adhikari and Busch, 2023) have been proposed. These protocols have shown promising enhancements in the transaction throughput of blockchain by processing non-conflicting transactions in parallel in multiple shards. However, none of these protocols have specifically explored the batch scheduling of transactions within a sharding environment, and they do not provide an extensive theoretical analysis for transaction scheduling.
@@
To process transactions parallelly in the sharding model, some research work (Kokoris-Kogias et al., 2018; Hellings and Sadoghi, 2021; Set and Park, 2022) has used two-phase locking for concurrency control. However, locks are expensive because when one process locks shared data for reading/writing, all other processes attempting to access the same data set are blocked until the lock is released, which lowers system throughput. Our scheduling algorithm does not use locks, as concurrency control is managed by the scheduler by processing non-conflicting transactions concurrently. In (Adhikari and Busch, 2023) the authors propose lockless blockchain sharding using multi-version concurrency control. However, they lack an analysis and technique for optimal scheduling. Moreover, they do not provide the benefits of locality and optimization techniques for transaction scheduling.
@@
Several works have been conducted on transaction scheduling in shared memory multi-core systems, distributed systems, and transactional memory. In a recent work (Busch et al., 2023), a stable scheduling algorithm was proposed specifically for software transactional memory systems, considering transactions generated by an adversarial model. Similarly, in research works (Attiya et al., 2015; Sharma and Busch, 2014, 2015), authors explored transaction scheduling in distributed transactional memory systems aimed at achieving better performance bounds with low communication costs. Authors in work (Busch et al., 2017) provided offline scheduling for transactional memory, where each transaction attempts to access an object, and once it obtains the object, it executes the transaction. In another (Busch et al., 2022), the author extended their work from offline to online scheduling for the transactional memory model. However, these works do not address transaction scheduling problems in the context of blockchain sharding. This is because, in the transactional memory model, the object is mobile, and once the transaction obtains the object, it immediately executes the transaction. In contrast, in blockchain sharding, the object is static in the shard, and there is a confirmation scheme to confirm and commit each subtransaction consistently in the respective accessing shard.
@@
In a recent work (Adhikari et al., 2024a, b), the authors provide a stability analysis of blockchain sharding considering adversarial transaction generation. However, their solution does not consider different types of sharding graphs, and they also do not provide a theoretical analysis of the optimal approximation for the scheduling algorithm.
3. Technical Preliminaries

### v3

@@
Abstract
Sharding is a promising technique for addressing the scalability issues of blockchain, and this technique is especially important for IoT, edge, or mobile computing. It divides the $n$ participating nodes into $s$ disjoint groups called shards, where each shard processes transactions in parallel. We examine batch scheduling problems on the shard graph $G_{s}$, where we find efficient schedules for a set of transactions. First, we present a centralized scheduler where one of the shards is considered as a leader, who receives the transaction information from all of the other shards and determines the schedule to process the transactions. For general graphs, where a transaction and its accessing objects are arbitrarily far from each other with a maximum distance $d$, the centralized scheduler provides $O(kd)$ approximation to the optimal schedule, where $k$ is the maximum number of shards each transaction accesses. Consequently, for a Clique graph where shards are at a unit distance from each other, we obtain $O(k)$ approximation to the optimal schedule. Next, we provide a centralized scheduler with a bucketing approach that offers improved bounds for the case where $G_{s}$ is a line graph, or the $k$ objects are randomly selected. Finally, we provide a distributed scheduler where shards do not require global transaction information. We achieve this by using a hierarchical clustering of the shards and using the centralized scheduler in each cluster. We show that the distributed scheduler has a competitive ratio of $O(\mathcal{A_{\mathcal{CS}}}\cdot\log d\cdot\log s)$, where $\mathcal{A_{\mathcal{CS}}}$ is the approximation ratio of the centralized scheduler. To our knowledge, we are the first to give provably fast transaction scheduling algorithms for blockchain sharding systems. We also present simulation results for our schedulers and compare their performance with a lock-based approach. The results show that our schedulers are generally better with up to 3x lower latency and 2x higher throughput.
Index Terms:
@@ Abstract
I Introduction
@@
Blockchains are known for their special features such as fault tolerance, transparency, non-repudiation, immutability, and security [1]. Thus, blockchains have been used in various domains, such as cryptocurrency [2, 3], healthcare [4], insurance [5], digital forensics [6, 7], supply chain management [8], edge and scientific computing [9]. Blockchain technology is also gaining interest from Transportation Systems[10], Mobile Edge Computing[11], cloud-edge-end and cooperative network [12], because of its secure, decentralized, and fault tolerance attributes [10, 11, 12]. However, the primary drawback of blockchain is its scalability and its performance, which is dependent on the throughput of the transactions [13, 14]. Thus, there is a research gap to address the scalability issues of blockchain.
@@
To append a new block in the blockchain network, each participating node needs to reach a consensus, which is a time and energy-consuming process [15, 16]. Moreover, each node is required to process and store all transactions, which leads to scalability issues in the blockchain system. To improve scalability and performance, sharding protocols have been proposed, such as Elastico [17], OmniLedger [18], RapidChain [19], Sharper [20], X-Shard[21] and ByShard [22]. However, these sharding protocols do not provide formal analysis for the schedule time complexity (i.e. how fast the transactions are processed). Sharding divides participating nodes into smaller groups called shards that allow the parallel processing of transactions. In the sharded blockchain, independent transactions are processed and committed in multiple shards concurrently, which improves the blockchain system’s throughput.
Figure 1: Edge, IoT and Mobile devices with blockchain sharding.
Mobile-edge computing data [11] and cloud-edge data [12] can be processed with a blockchain that utilizes sharding, which reduces storage, communication, and computation costs [19]. Figure 1 shows sharding with IoT and mobile devices, where each shard operates on edge computing servers, connects with IoT devices, and handles transaction execution. In Figure 1, Shards $1$ and $2$ are responsible for processing and handling mobile data, Shard $3$ stores data related to transportation and connected with devices such as airlines, cars, and trains, and Shard $4$ stores data related to IoT devices, connecting with devices like smartwatches, Wi-Fi routers, and banks. Transactions can be generated in any shard, and the shard is responsible for processing a particular transaction. We assume that these devices generate transactions and send them to the shards for storing and processing. Our goal is to process/store these transactions as fast as possible. Our scheduler determines the order in which these transactions are processed in a conflict-free manner, as some transactions might conflict by attempting to store or update the same object (e.g., entity or account).
@@
We consider a blockchain system consisting of $n$ nodes, which are further divided into $s$ shards, and each shard consists of a set of nodes. Shards are connected in a graph network $G_{s}$ with a diameter $D$, and each shard holds a subset of the objects (accounts). We assume that transactions are distributed across the shards, and each transaction accesses at most $k$ accounts. A transaction $T$ initially is in one of the shards, which is called the home shard for $T$. Similar to other sharding systems [22, 15, 16], each transaction $T$ is split into subtransactions, where each subtransaction accesses an account. A subtransaction of $T$ is sent to the destination shard that holds the respective account. The maximum distance between the home shard of a transaction and the respective destination shards in $G_{s}$ is at most $d\leq D$. In other words, the transaction and its accessing shards (objects) are arbitrarily far from each other with a maximum distance of $d$.
All home shards process transactions concurrently. A problem occurs when conflicting transactions try to access the same account simultaneously. In such a case, the conflict prohibits the transactions to be committed concurrently and forces them to serialize [16]. Our proposed scheduling algorithms coordinate the home shards and destination shards to process the transactions (and respective subtransactions) in a conflict-free manner. Each destination shard maintains a local blockchain of the subtransactions that are sent to it. The global blockchain can be constructed (if needed) by combining the local blockchains at the shards [15].
We consider batch problem instances where a set of transactions ${\mathcal{T}}$ are generated by (IoT and mobile) devices that reside at shards. We study the shard model in different graph topologies such as General graph, Clique (Complete) graph, and Line graphs and provide the analysis for execution time. Considering the different graph topologies for the IoT and mobile computing is important as presented in[23, 24].
@@
The scheduler determines the time step for each transaction $T_{i}\in{\mathcal{T}}$ to process and commit. The execution of our scheduling algorithm is synchronous, where the timeline is divided into time steps referred to as rounds. Similar to the work in [16], we consider that the duration of a round is sufficient to allow the execution of the PBFT consensus algorithm [25] in each shard. A round is also the time to send a message between shards in a unit distance. The main goal of a scheduling algorithm is to efficiently and fairly process all transactions while minimizing the total execution time (makespan) until all transactions are either committed or aborted. Unlike previous sharding approaches [22, 18, 19], our scheduling algorithms do not require locking mechanisms for concurrency control. Hence, we can also say our algorithms are lock-free.
Shards are Connected as Approximation to Optimal Schedule of
@@
General Graph $O(kd)$ $O(kd\cdot\log d\cdot\log s)$
Clique Graph with unit distance $O(k)$ $O(k)$ [same with centralized]
Hypercube, Butterfly, and $g$-dimensional Grid Graph $O(k\log s)$ $O(k\cdot\log d\cdot\log^{2}s)$
General Graph where transaction access $k$ objects randomly $O(k(k+\log s)\log D)$ $O(k(k+\log s)\log D\cdot\log d\cdot\log s)$
Line Graph $O(k\sqrt{d}\log D)$ $O(k\sqrt{d}\log D\cdot\log d\cdot\log s)$
@@
TABLE I: Summary of the proposed centralized and distributed scheduler approximation to the optimal schedule. The used notations are as follows: $s$ represents the total number of shards, $k$ denotes the maximum number of shards (objects) accessed by each transaction, $d$ denotes the worst distance between any transaction (home shard) and its accessing objects (destination shard), $D$ denotes the diameter of shard graph $G_{s}$.
I-A Contributions
@@
To our knowledge, we give the first provably efficient batch transaction scheduling algorithms for blockchain sharding systems. Table I has a summary of our contributions which we also described as follows:
@@
  • Centralized Scheduler: First, we provide a Centralized scheduler where one of the shards is considered as a leader, and all of the other shards send their transaction information to the leader, and the leader determines the schedule to process the transactions. Consider a general graph $G_{s}$ with diameter $D$. Suppose that each transaction accesses at most $k$ objects, where the distance between a transaction and its accessing objects is at most $d\leq D$. The centralized scheduler provides an $O(kd)$ approximation to the optimal schedule. Hence, when $G_{s}$ is a Clique graph, where shards are at a unit distance from each other (thus, $d=1$), we obtain an $O(k)$ approximation to the optimal schedule. Moreover, we obtain an $O(k\log s)$ approximation for the Hypercube, Butterfly, and $g$-dimensional Grid where $g=O(\log s)$, since in all these cases $d=O(\log s)$.
@@
  • Centralized Scheduler with Buckets: Next, we give the centralized scheduler with a bucketing approach. We divide the transactions into buckets $B_{0},B_{1},\ldots$, according to the maximum distance between the home shard of the transaction and its accessing objects (i.e., destination shards). $B_{i}$ consists of the transactions whose maximum distance between the home shard and its destination shard is in range $[2^{i},2^{i+1})$. This approach offers improved bounds for special cases: for the General graph where the $k$ objects are chosen randomly per transaction we give $O(k(k+\log s)\log D)$ approximation, and for the Line graph, we provide $O(k\sqrt{d}\log D)$ approximation to the optimal schedule.
@@
  • Distributed Scheduler: Finally, we give a distributed scheduler where shards are not aware of global transaction information. We use a hierarchical clustering approach [26] for the shard network so that independent transactions in different clusters can be scheduled and committed concurrently. The basic idea is to invoke the centralized scheduler within each cluster of the hierarchy to make a distributed scheduler. We show that the distributed scheduler has a competitive ratio of $O(\mathcal{A_{\mathcal{CS}}}\cdot\log d\cdot\log s)$, where $\mathcal{A_{\mathcal{CS}}}$ is the approximation ratio of the centralized scheduler.
  • Simulation Results: We simulate our proposed transaction scheduling algorithms and analyze their average throughput, latency, and number of message exchanges. The simulation results show that our schedulers are better than a lock-based (Byshard [22]) approach by achieving up to $2$x higher throughput, $3$x lower latency, and $6$x fewer message exchanges.
@@
Paper Organization: The rest of this paper is structured as follows: Section II provides related works. Section III describes the preliminaries for this study and the sharding model. Section IV presents a centralized scheduling algorithm. In Section V, we provide a centralized scheduler with the bucketing approach. Section VI generalizes the techniques to a fully distributed setting. In Section VII, we provide the simulation results of each proposed algorithm. Finally, we give our conclusions in Section VIII.
II Related Work
@@
Extensive research has been proposed to address the scalability issues of blockchain in the consensus layer [27, 28, 29, 30, 31, 32, 33]. While these proposed algorithms and protocols have made some improvements in scalability, the system’s performance still degrades as the network size increases. To tackle the scalability issue of blockchain, various sharding protocols [17, 18, 19, 22, 34, 20, 15, 35] have been proposed. These protocols have shown promising enhancements in the transaction throughput of blockchain by processing non-conflicting transactions in parallel in multiple shards. However, none of these protocols have specifically explored the batch scheduling of transactions within a sharding environment, and they do not provide an extensive theoretical analysis for transaction scheduling. There are many research work has been done to securely store and process the IoT and mobile computing data in blockchain[36, 37, 11], and blockchain sharding [38, 39]. However, these works also lack the theoretical analysis for the optimal schedule.
@@
To process transactions parallelly in the sharding model, some research work [18, 22, 40] has used two-phase locking for concurrency control. However, locks are expensive because when one process locks shared data for reading/writing, all other processes attempting to access the same data set are blocked until the lock is released, which lowers system throughput. Moreover, locks if not handled and released properly may cause deadlocks. Our scheduling algorithm does not use locks, as concurrency control is managed by the scheduler by processing non-conflicting transactions concurrently. In [15] the authors propose lockless blockchain sharding using multi-version concurrency control. However, they lack an analysis and technique for optimal scheduling. Moreover, they do not provide the benefits of locality and optimization techniques for transaction scheduling.
@@
Several works have been conducted on transaction scheduling in shared memory multi-core systems, distributed systems, and transactional memory. In a recent work [41], a stable scheduling algorithm was proposed specifically for software transactional memory systems, considering transactions generated by an adversarial model. Similarly, in research works [42, 43, 44], authors explored transaction scheduling in distributed transactional memory systems aimed at achieving better performance bounds with low communication costs. Authors in work [45] provided offline scheduling for transactional memory, where each transaction attempts to access an object, and once it obtains the object, it executes the transaction. In another work [46], the authors extended their analysis from offline to online scheduling for the transactional memory model. However, these works do not address transaction scheduling problems in the context of blockchain sharding. This is because, in the transactional memory model, the object is mobile, and once the transaction obtains the object, it immediately executes the transaction. In contrast, in blockchain sharding, the object is static in the shard, and there is a confirmation scheme to confirm and commit each subtransaction consistently in the respective accessing shard.
@@
In a recent work [16], the authors provide a stability analysis of blockchain sharding considering adversarial transaction generation. However, their solution does not consider different types of sharding graphs, and they also do not provide a theoretical analysis of the optimal approximation for the scheduling algorithm.
III Technical Preliminaries